Problem 12-2 Trading securities; bond investment; effective interest [LO12-1, 12-3]

Fuzzy Monkey Technologies, Inc., purchased as a short-term investment $250 million of 8% bonds, dated January 1, on January 1, 2018. Management intends to include the investment in a short-term, active trading portfolio. For bonds of similar risk and maturity the market yield was 10%. The price paid for the bonds was $228 million. Interest is received semiannually on June 30 and December 31. Due to changing market conditions, the fair value of the bonds at December 31, 2018, was $240 million.

Required:
1. to 3. Prepare the relevant journal entries on the respective dates (record the interest at the effective rate).
4-a. At what amount will Fuzzy Monkey report its investment in the December 31, 2018, balance sheet?
4-b. Prepare any entry necessary to achieve this reporting objective.
5. How would Fuzzy Monkey's 2018 statement of cash flows be affected by this investment?

How would Fuzzy Monkey's 2018 statement of cash flows be affected by this investment? (Do not round intermediate calculations. Enter your answers in millions rounded to 1 decimal place, (i.e., 5,500,000 should be entered as 5.5).)

Operating cash flow $20.0selected answer incorrect million outflowselected answer correct
Investing cash flow $0.0selected answer correct million no effectselected answer correct
Answer #1

Journal entry from 1 to 3: (in millions)

Date Account title Debit    Credit
Jan 1, 2018 Investment in 8% bonds $250
Discount on bond investment (250 - 228) $22
Cash $228
Jun 30, 2018 Cash ($250 * 8/2 %) $10
Discount on bond investment $1.4
Interest revenue ($228 * 10/2 %) $11.4
Dec 31, 2018 Cash ($250 * 8/2 %) $10
Discount on bond investment $1.5
Interest revenue [($228 +1.4) * 10/2 %] $11.5

Answer to part 4 a:

The company will report it's investment at $240 million in balance sheet on 31, december 2018. (as mentioned in the question)

Answer to part 4 b:

Journal entry to achieve the above adjustment:(IN MILLIONS)

Account title Debit Credit
Fair value adjustment $9.1
Unrealized holding gain - NI $9.1

WORKING NOTE:

investment in bonds = $250 million

discounts on bonds on 1/1 = $22 million

amortisation value = $1.4 million and $1.5 million = $2.9 million

Balance of bonds on 31/12 = 250 -22 + 2.9 = $230. 9 million

Fair value as on 31/12 = $240 million

Fair value adjustment = 240 - 230.9 = $9.1 million

Answer to part 5:

operating cash outflow = $228 - $10 - $0 = $208 millions

Kindly leave feedback !Thank you!